# SyncDiff: Synchronized Motion Diffusion for Multi-Body Human-Object Interaction Synthesis

### v4

@@ 1 Introduction
In practice, we often observe high-frequency, small-amplitude vibrations with semantics be overshadowed by low-frequency, large-scale movements (e.g., A brush only contacts the teapot, without relative frictions between them). To solve this, we propose frequency decomposition, which factorizes all individual and relative motions into low and high frequency components, and supervises high-frequency ones in the frequency domain independently.
@@ 3 Method
Figure 2: Overview of SyncDiff. The light blue boxes show the inference process with explicit synchronization steps performed every $s$ step. For denoising steps irrelevant to explicit synchronization (those marked as “$(s-1)$ times”), the noise level is set to $\sigma_{t}$. For the calculation from $\hat{x}$ to $\hat{\mu}$, no noise is added. For the calculation of $\hat{x}_{t-1}$ based on $\hat{x}_{t}$ and $\hat{\mu}_{t}$ in explicit synchronization step, the noise level is $\sigma_{t}^{{}^{\prime}}$. Please refer to Section 3.6 for more details. The light green box illustrates the architecture of denoising backbone (Section 3.4).
@@ 3.1 Problem Definition
Consider an interaction scenario comprising $n$ articulated skeletons $h_{1},h_{2},\dots,h_{n}$ and $m$ rigid bodies $o_{1},o_{2},\dots,o_{m}$, which we uniformly define as “body”. In our experiments, these skeletons are MANO [68] hands or SMPL-X [57] humans, whose motions can be reconstructed with joint information and intrinsic shape parameter $\beta$. Details can be found in Section  C.2. Our task is to synthesize body motions with known action labels, object categories and geometries, and shape parameters $\beta_{i\in[1,n]}$ of each skeleton. We highlight that our task releases the needs of either predefined object motions [42] or HOI key frames [13] that are used in other existing works.
@@ 3.2 Motion Representation
For two bodies $\textbf{b}_{1}$ and $\textbf{b}_{2}$, we use $x_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}}$ to denote the relative motion of $\textbf{b}_{2}$ in $\textbf{b}_{1}$’s coordinate system. We require that $\textbf{b}_{1}$ is a rigid body, and omit relative representation between articulated skeletons. If $\textbf{b}_{2}$ is also a rigid body, then
$x_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}}=[\textbf{t}_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}},\textbf{q}_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}}]=[\textbf{q}_{\textbf{b}_{1}}^{-1}(\textbf{t}_{\textbf{b}_{2}}-\textbf{t}_{\textbf{b}_{1}}),\textbf{q}_{\textbf{b}_{1}}^{-1}\textbf{q}_{\textbf{b}_{2}}].$ (1)
If $\textbf{b}_{2}$ is an articulated skeleton, the position of one of its joints p under individual representation will become $\textbf{p}^{{}^{\prime}}$ in $\textbf{b}_{1}$’s coordinate system, where
$\textbf{p}^{{}^{\prime}}=\textbf{q}_{\textbf{b}_{1}}^{-1}(\textbf{p}-\textbf{t}_{\textbf{b}_{1}}).$ (2)
After deriving all individual and relative motion representations, we concatenate all of them together into a high-order representation $x$, including $\{x_{o_{j\in[1,m]}}\}$, $\{x_{h_{i\in[1,n]}}\}$, $\{x_{o_{j_{2}}\rightarrow o_{j_{1}}}\mid j_{1},j_{2}\in[1,m],j_{1}\neq j_{2}\}$, $\{x_{h_{i}\rightarrow o_{j}}\mid i\in[1,n],j\in[1,m]\}$. Thus, $x\in\mathbb{R}^{N\times D_{\text{sum}}}$, where
$D_{\text{sum}}$ $=7m+3Dn+7m(m-1)+3Dmn$
$=7m^{2}+3Dn(m+1).$ (3)
Although $x$ is composed of individual and relative representations, the final synthesized motions only involve individual motions, and those relative ones serve merely as auxiliary components.
@@ 3.3 Frequency Decomposition
As body number increases, action semantics become more dependent on subtle, high-frequency interactions between bodies (e.g., periodic frictions between the teapot and brush during the action of brushing). Unfortunately, using existing works to denoise based on time-domain trajectories often results in these high-frequency components being overshadowed by common low-frequency, large-scale patterns (e.g., objects only move and contact each other, without relative interactions). To better model these high-frequency components, inspired by GID [43], we explicitly represent them in the frequency domain and supervise them independently in the loss terms described in Section  3.5.
@@
where $a_{l},b_{l}\in\mathbb{R}^{D_{\text{sum}}}$ are the coefficients computed by FFT. Note that here $u$ denotes the frame id, which is different from the noise timestep $t$ for the diffusion model later. To prevent networks from overfitting high-frequency noises in mocap datasets, we select a cutoff boundary $L\in[4,\frac{N}{4})$, and directly discard signals with frequencies higher than $\phi_{L}/2\pi$. In practice, we select $L=16$ (Reasons can be found in Section  B.1. We then divide the remaining signals into low-frequency components ($x_{\text{dc}}$) and high-frequency components ($x_{\text{ac}}$):
@@ 3.4 Model Architecture
We jointly model all individual and relative motions in a high-order representation with one single diffusion model (Figure 2). The latent diffusion [67] paradigm is adopted, where action and object label features are extracted from pretrained CLIP  [64], and object geometry features are encoded by BPS  [59]. To facilitate batch operations, we pad all trajectories into the same length, and use 0/1 padding masks to indicate the positions that need to be generated. Concatenate together label/geometry features, padding masks, noise timestep embeddings, and shape parameters $\beta_{i\in[1,n]}$ to form the condition vector. The condition vector is replicated and combined with $x_{t,\text{dc}}$ and $x_{t,\text{F}}$ decomposed from noised $x_{t}$, respectively. They are then projected into the latent space to be denoised by a transformer-based backbone. Let the predicted results be $\widehat{x_{\text{dc}}}$ and $\widehat{x_{\text{F}}}$, the latter of which is reconstructed as $\widehat{x_{\text{ac}}}$. The final denoised motion sequence is recomposed by $\widehat{x}=\widehat{x_{\text{dc}}}+\widehat{x_{\text{ac}}}$. Please refer to Section  D.2,  D.3 for hyperparameters of model architecture and training/inference processes.
@@ 3.5 Loss Functions
For data sample scores, our method is similar to the standard reconstruction loss, except that we supervise $x_{\text{dc}}$ and $x_{\text{ac}}$ separately. We denote $\mathcal{L}_{\text{dc}}=\|\widehat{x_{\text{dc}}}-x_{\text{dc}}\|_{2}^{2}$, $\mathcal{L}_{\text{ac}}=\|\widehat{x_{\text{ac}}}-x_{\text{ac}}\|_{2}^{2}$ and $\mathcal{L}_{\text{norm}}=\sum\limits_{j=1}^{m}\|1-|\hat{\textbf{q}}_{j}|\|_{2}^{2}$. The last one is used to induce the norms of the quaternions representing rigid body rotations to be as close to $1$, where $\hat{x}_{o_{j}}=[\hat{\textbf{t}}_{j},\hat{\textbf{q}}_{j}]$.
@@
Proof details of the equivalence between $\mathcal{L}_{\text{align}}$ and alignment scores are provided in Section  A.2.
@@ 3.6 Explicit Synchronization in Inference Time
Relying solely on alignment losses only indirectly enhances synchronization. To directly improve the synthesis quality and diversity, as well as provide stronger theoretical guarantees, our strategy is to introduce an explicit synchronization process, which is mathematically equivalent to maximum total likelihood sampling during inference time, aiming to leverage both data sample scores and alignment scores to address this problem. Since the synchronization step is time-consuming, to balance performance and efficiency, we perform synchronization operations every $s(s\ll T)$ steps, where $T$ is the total number of denoising steps, as is shown in Figure 2. In practice, we take $s=50$, $T=1000$, the reason for which is shown in Section  B.2. For the predicted motion $\hat{x}_{t}$ at step $t\in[1,T]$, according to the diffusion formula [30], without synchronization, the next step would be:
@@
where $\text{comb}(x_{\textbf{a}},x_{\textbf{b}\rightarrow\textbf{a}})$ utilizes the individual motion of a and relative motion between b and a to compute b’s motion. Its formula is in Section  C.1. $\overline{\lambda}=\frac{\lambda_{\text{exp}}}{R}\sum\limits_{r=1}^{R}\frac{1}{2\sigma_{t_{r}}^{2}}$, where $R=T/s=20$ is the number of synchronization steps, $t_{1},t_{2},\dots,t_{R}$ are the corresponding timesteps, and $\sigma_{t_{1}}$, $\sigma_{t_{2}}$, $\dots$, $\sigma_{t_{R}}$ are the original correspondent noise scales (without synchronization). Finally, $\sigma^{{}^{\prime}}=\sqrt{\frac{\sigma^{2}}{1+2\sigma^{2}\overline{\lambda}}}$. Specifically, when $m=1$, we do not perform synchronization for this part, and the denoising formula is identical to that without synchronization.
@@
Definitions of $\overline{\lambda}$ and $\sigma^{{}^{\prime}}$ are the same as Eq. 9. Function rel is defined in Section  3.5.
After synchronization operations and deriving $\hat{x}^{{}^{\prime}}$, it is again used for further denoising. Between two adjacent synchronization steps, we still directly use Eq. 8 for stepwise denoising, and do not perform synchronization operations. We demonstrate in Section  A.2, that the above equations are equivalent to maximum likelihood sampling from the newly computed Gaussian distribution based on both data sample scores and alignment scores. To help readers understand the inference process with explicit synchronization steps more intuitively, we also provide the pseudocode in Section  A.3.
@@ 4 Experiments
In this section, we first introduce some basic experimental settings (Section  4.1,  4.2), and then demonstrate the visual and quantitative results in comparison to several baselines (Section  4.3) and ablation settings (Section  4.4). We also provide in-depth analyses for the results.
@@ 4.1 Datasets
To examine our method’s generalizability across various multi-body interaction configurations, we utilize five datasets with different interaction scenarios: TACO [48] (two hands and two objects), CORE4D [109] (two people and one object), GRAB [81] (one or two hands and one object), OAKINK2 [108] (two hands and one to three objects), and BEHAVE [2] (one human and one object). We describe data splits for each dataset below. Detailed dataset statistics and split sizes are shown in Section  D.1.
@@ 4.2 Evaluation Metrics
(1) Contact-based metrics measure the contact plausibility of hand-object/human-object interactions and the extent of motion coordination among different bodies. We use the Contact Surface Ratio (CSR) for hand-object settings and the Contact Root Ratio (CRR) for human-object settings to denote the proportion of frames where hand-object/human-object contact occurs. Contact is defined as the hand mesh being within $5$mm of at least one object for CSR, and the two hand roots of a human consistently being within $3$cm of at least one object for CRR. When there are multiple hands or humans, we take the average among all of them. We label the frames based on whether hand-object contact occurs for ground-truth or synthesized motions, and then compute their Intersection-over-Union (IoU), denoted as CSIoU. To eliminate potential ambiguities, we’ve provided corresponding pseudocodes in Section  C.5. Besides, Interpenetration Volume (IV) and Interpenetration Depth (ID) are incorporated to assess penetration between different bodies.
(2) Motion semantics metrics evaluate high-level motion semantics and its distributions, comprising Recognition Accuracy (RA), Fréchet Inception Distance (FID), Sample Diversity (SD), and Overall Diversity (OD). Following existing evaluations for motion synthesis [42, 41], we train a network to extract motion features and predict action labels using ground-truth motion data. For better feature semantics, the network is trained on the combination of all train, val, and test splits. RA denotes the action recognition accuracy of the network on synthesized motions. FID measures the difference in feature distributions of generated and ground-truth motions. Following DiffH2[13], SD represents the mean Euclidean distance between multiple generated wrist trajectories in a single sample, and OD refers to mean distance between all generated trajectories in a dataset split. To measure semantics from a perceptual aspect, we conduct user studies, comparing our methods with baselines, which is shown in Section  4.5.
@@ 4.3 Comparison to Existing Methods
In Figure  3, our method features more realistic contacts and more stable grasping. As is shown in Table 1, 2, and 5, our method outperforms MACS and DiffH2O by a large margin with better CSIoU, CSR, IV, and ID. The reason is that our method features more robust alignment and synchronization mechanisms to ensure synchronization which are excluded in existing methods. In the results from MACS and DiffH2O in Figure  4, the brush often penetrates the plate, and the pose of the brush does not guarantee the bristles being pressed closely against the plate, effectively completing the action. FID and RA further indicate that the motions generated by our method are more semantically realistic. This is caused by the more precise object-object interactions in SyncDiff, and our separation of motions at different frequencies ensures that subtle high-frequency periodic movements are not overshadowed, which are crucial for identifying action types. Figure  5 demonstrates the benefits of our synchronization mechanism. Although MACS uses relative representations for some hand-object pairs to ensure firm grasp, due to the lack of synchronization on the complete graphical model, conflicts arise between the left hand and the knife, which should not directly interact. Figure 6 indicates that with a higher demand for coordination quality among objects, our method can also address hand-object and object-object synchronization effectively.
@@
As shown in Table  3,  4 and Figure 7,  8, our method outperforms CG-HOI and OMOMO in contact-based and semantics metrics and obtains the most realistic qualities. Synthesized motions from existing methods suffer from unnatural grasping poses and arm-object interpenetration, while our method mitigates these issues.

### v5

@@ 1 Introduction
In practice, we often observe high-frequency, small-amplitude vibrations with semantics be overshadowed by low-frequency, large-scale movements (e.g., A brush only contacts the teapot without effective frictions between them). To solve this, we propose frequency decomposition, which factorizes all individual and relative motions into low and high frequency components, and supervises them in the time domain and frequency domain separately.
@@ 3 Method
[…]
@@ 3.1 Problem Definition
Consider an interaction scenario comprising $n$ articulated skeletons $h_{1},h_{2},\dots,h_{n}$ and $m$ rigid bodies $o_{1},o_{2},\dots,o_{m}$, which we uniformly define as “body”. In our experiments, these skeletons are MANO [68] hands or SMPL-X [57] humans, whose motions can be reconstructed with joint information and intrinsic shape parameter $\beta$. Details can be found in Section B.2. Our task is to synthesize body motions with known action labels, object categories and geometries, and shape parameters $\beta_{i\in[1,n]}$ of each skeleton. We highlight that our task releases the needs of either predefined object motions [42] or HOI key frames [13] that are used in other existing works.
@@ 3.2 Motion Representation
For two bodies $\textbf{b}_{1}$ and $\textbf{b}_{2}$, we use $x_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}}$ to denote the relative motion of $\textbf{b}_{2}$ in $\textbf{b}_{1}$’s coordinate system. We require that $\textbf{b}_{1}$ is a rigid body, and omit relative representation between articulated skeletons. If $\textbf{b}_{2}$ is also a rigid body, then $x_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}}=[\textbf{t}_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}},\textbf{q}_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}}]$, where $\textbf{t}_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}}=\textbf{q}_{\textbf{b}_{1}}^{-1}(\textbf{t}_{\textbf{b}_{2}}-\textbf{t}_{\textbf{b}_{1}})$, and $\textbf{q}_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}}=\textbf{q}_{\textbf{b}_{1}}^{-1}\textbf{q}_{\textbf{b}_{2}}$. If $\textbf{b}_{2}$ is an articulated skeleton, the position of one of its joints p under individual representation will become $\textbf{p}^{{}^{\prime}}$ in $\textbf{b}_{1}$’s coordinate system, where $\textbf{p}^{{}^{\prime}}=\textbf{q}_{\textbf{b}_{1}}^{-1}(\textbf{p}-\textbf{t}_{\textbf{b}_{1}})$.
After deriving all individual and relative motion representations, we concatenate all of them together into a high-order representation $x$, including $\{x_{o_{j\in[1,m]}}\}$, $\{x_{h_{i\in[1,n]}}\}$, $\{x_{o_{j_{2}}\rightarrow o_{j_{1}}}\mid j_{1},j_{2}\in[1,m],j_{1}\neq j_{2}\}$, $\{x_{h_{i}\rightarrow o_{j}}\mid i\in[1,n],j\in[1,m]\}$. Thus, $x\in\mathbb{R}^{N\times D_{\text{sum}}}$, where $D_{\text{sum}}=7m+3Dn+7m(m-1)+3Dmn$. Although $x$ is composed of individual and relative representations, the final synthesized motions only involve individual motions, and those relative ones serve merely as auxiliary components.
@@ 3.3 Frequency Decomposition
As body number increases, action semantics become more dependent on subtle, high-frequency interactions between bodies (e.g., periodic frictions between the teapot and brush during the action of brushing). Unfortunately, using existing works to denoise based on time-domain trajectories often results in these high-frequency components being overshadowed by common low-frequency, large-scale patterns (e.g., objects only move and contact each other, without relative interactions). To better model these high-frequency components, inspired by GID [43], we explicitly represent them in the frequency domain and supervise them independently in the loss terms described in Section 3.5.
@@
where $a_{l},b_{l}\in\mathbb{R}^{D_{\text{sum}}}$ are the coefficients computed by FFT. Note that here $u$ denotes the frame id, which is different from the noise timestep $t$ for the diffusion model later. To prevent networks from overfitting high-frequency noises in mocap datasets, we select a cutoff boundary $L\in[4,\frac{N}{4})$, and directly discard signals with frequencies higher than $\phi_{L}/2\pi$. In practice, we select $L=16$, which strikes a balance between motion fidelity conservation and simplicity (See Section B.1). We then divide the remaining signals into low-frequency components ($x_{\text{dc}}$) and high-frequency components ($x_{\text{ac}}$):
@@ 3.4 Model Architecture
We jointly model all individual and relative motions in a high-order representation with one single diffusion model (Figure 2). The latent diffusion [67] paradigm is adopted, where action and object label features are extracted from pretrained CLIP  [64], and object geometry features are encoded by BPS  [59]. To facilitate batch operations, we pad all trajectories into the same length, and use 0/1 padding masks to indicate the positions that need to be generated. Concatenate together label/geometry features, padding masks, noise timestep embeddings, and shape parameters $\beta_{i\in[1,n]}$ to form the condition vector. The condition vector is replicated and combined with $x_{t,\text{dc}}$ and $x_{t,\text{F}}$ decomposed from noised $x_{t}$, respectively. They are then projected into the latent space to be denoised by a transformer-based backbone. Let the predicted results be $\widehat{x_{\text{dc}}}$ and $\widehat{x_{\text{F}}}$, the latter of which is reconstructed as $\widehat{x_{\text{ac}}}$. The final denoised motion sequence is recomposed by $\widehat{x}=\widehat{x_{\text{dc}}}+\widehat{x_{\text{ac}}}$. Please refer to Section D.2D.3 for model hyperparameters.
@@ 3.5 Loss Functions
For data sample scores, our method is similar to the standard reconstruction loss, except that we supervise $x_{\text{dc}}$ and $x_{\text{ac}}$ separately. We denote $\mathcal{L}_{\text{dc}}=\|\widehat{x_{\text{dc}}}-x_{\text{dc}}\|_{2}^{2}$, $\mathcal{L}_{\text{ac}}=\|\widehat{x_{\text{ac}}}-x_{\text{ac}}\|_{2}^{2}$ and $\mathcal{L}_{\text{norm}}=\sum\limits_{j=1}^{m}\|1-|\hat{\textbf{q}}_{j}|\|_{2}^{2}$. The last one is used to induce the norms of the quaternions representing rigid body rotations to be as close to 111, where $\hat{x}_{o_{j}}=[\hat{\textbf{t}}_{j},\hat{\textbf{q}}_{j}]$.
@@
Proof details of the equivalence between $\mathcal{L}_{\text{align}}$ and alignment scores are provided in Section A.2.
@@ 3.6 Explicit Synchronization in Inference Time
Relying solely on alignment losses only indirectly enhances synchronization. To directly improve synthesis quality and diversity, as well as provide stronger theoretical guarantees, we introduce an explicit synchronization process, which is mathematically equivalent to maximum total likelihood sampling during inference time, aiming to leverage both data sample scores and alignment scores to address this problem. Since the synchronization step is time-consuming, to balance performance and efficiency, we perform synchronization operations every $s(s\ll T)$ steps, where $T$ is the total number of denoising steps, as is shown in Figure 2. In practice, we take $s=50$, $T=1000$, to ensure synchronization while improving inference speed, as further detailed in Section B.2. For the predicted motion $\hat{x}_{t}$ at step $t\in[1,T]$, according to the diffusion formula [30], without synchronization, the next step would be:
@@
where $\text{comb}(x_{\textbf{a}},x_{\textbf{b}\rightarrow\textbf{a}})$ utilizes the individual motion of a and relative motion between b and a to compute b’s motion. Its formula is in Section C.1. $\overline{\lambda}=\frac{\lambda_{\text{exp}}}{R}\sum\limits_{r=1}^{R}\frac{1}{2\sigma_{t_{r}}^{2}}$, where $R=T/s=20$ is the number of synchronization steps, $t_{1},t_{2},\dots,t_{R}$ are the corresponding timesteps, and $\sigma_{t_{1}}$, $\sigma_{t_{2}}$, $\dots$, $\sigma_{t_{R}}$ are the original correspondent noise scales (without synchronization). Finally, $\sigma^{{}^{\prime}}=\sqrt{\frac{\sigma^{2}}{1+2\sigma^{2}\overline{\lambda}}}$. Specifically, when $m=1$, we do not perform synchronization for this part, and the denoising formula is identical to that without synchronization.
@@
Definitions of $\overline{\lambda}$ and $\sigma^{{}^{\prime}}$ are the same as Eq. 6. Function rel is defined in Section 3.5.
After synchronization operations and deriving $\hat{x}^{{}^{\prime}}$, it is again used for further denoising. Between two adjacent synchronization steps, we still directly use Eq. 5 for stepwise denoising, and do not perform synchronization operations. We demonstrate in Section A.2, that the above equations are equivalent to maximum likelihood sampling from the newly computed Gaussian distribution based on both data sample scores and alignment scores.
@@ 4 Experiments
In this section, we first introduce some basic experimental settings (Sections 4.14.2), and then demonstrate the visual and quantitative results in comparison to several baselines (Section 4.3) and ablation settings (Section 4.4). We also provide in-depth analyses for the results.
@@ 4.1 Datasets
To examine our method’s generalizability across various multi-body interaction configurations, we utilize five datasets with different interaction scenarios: TACO [48] (two hands and two objects), CORE4D [109] (two people and one object), GRAB [81] (one or two hands and one object), OAKINK2 [108] (two hands and one to three objects), and BEHAVE [2] (one human and one object). We describe data splits for each dataset below. Detailed dataset statistics and split sizes are shown in Section D.1.
@@ 4.2 Evaluation Metrics
(1) Contact-based metrics measure the contact plausibility of hand-object/human-object interactions and the extent of motion coordination among different bodies. We use the Contact Surface Ratio (CSR) for hand-object settings and the Contact Root Ratio (CRR) for human-object settings to denote the proportion of frames where hand-object/human-object contact occurs. Contact is defined as the hand mesh being within 555mm of at least one object for CSR, and the two hand roots of a human consistently being within 333cm of at least one object for CRR. When there are multiple hands or humans, we take the average among all of them. We label the frames based on whether hand-object contact occurs for ground-truth or synthesized motions, and then compute their Intersection-over-Union (IoU), denoted as CSIoU. To eliminate potential ambiguities, we’ve provided corresponding pseudocodes in Section C.5. Besides, Interpenetration Volume (IV) and Interpenetration Depth (ID) are incorporated to assess penetration between different bodies.
(2) Motion semantics metrics evaluate high-level motion semantics and its distributions, comprising Recognition Accuracy (RA), Fréchet Inception Distance (FID), Sample Diversity (SD), and Overall Diversity (OD). Following existing evaluations for motion synthesis [42, 41], we train a network to extract motion features and predict action labels using ground-truth motion data. For better feature semantics, the network is trained on the combination of all train, val, and test splits. RA denotes the action recognition accuracy of the network on synthesized motions. FID measures the difference in feature distributions of generated and ground-truth motions. Following DiffH2[13], SD represents the mean Euclidean distance between multiple generated wrist trajectories in a single sample, and OD refers to mean distance between all generated trajectories in a dataset split. To measure semantics from a perceptual aspect, we conduct user studies, comparing our methods with baselines, which is shown in Section 4.5.
@@ 4.3 Comparison to Existing Methods
In Figure 3, our method features more realistic contacts and more stable grasping. As is shown in Table 1, 2, and 5, our method outperforms MACS and DiffH2O by a large margin with better CSIoU, CSR, IV, and ID. The reason is that our method features more robust alignment and synchronization mechanisms to ensure synchronization which are excluded in existing methods. In the results from MACS and DiffH2O in Figure 4, the brush often penetrates the plate, and the pose of the brush does not guarantee the bristles being pressed closely against the plate, effectively completing the action. FID and RA further indicate that the motions generated by our method are more semantically realistic. This is caused by the more precise object-object interactions in SyncDiff, and our separation of motions at different frequencies ensures that subtle high-frequency periodic movements are not overshadowed, which are crucial for identifying action types. Figure 5 demonstrates the benefits of our synchronization mechanism. Although MACS uses relative representations for some hand-object pairs to ensure firm grasp, due to the lack of synchronization on the complete graphical model, conflicts arise between the left hand and the knife, which should not directly interact. Figure 6 indicates that with a higher demand for coordination quality among objects, our method can also address hand-object and object-object synchronization effectively.
@@
As shown in Tables 34 and Figure 78, our method outperforms CG-HOI and OMOMO in contact-based and semantics metrics and obtains the most realistic qualities. Synthesized motions from existing methods suffer from unnatural grasping poses and arm-object interpenetration, while our method mitigates these issues.